# Half-metallicity and wandering axis ferromagnetism in Mn substituted Fe2TiSn

### v1

@@ Abstract
We investigate the effect of Mn substitution in Fe2Ti1-xMnxSn on electronic structure, magnetic and electrical transport properties. The spin-polarized density of states calculations using density-functional theory (DFT), yields a half-metallic ground state in Mn-rich compositions. Localized magnetic moments at Mn sites that interact through the cloud of conduction electrons formed by Fe and Ti atoms are also predicted. Electrical resistivity and magneto-transport measurements reveal a Kondo-like ground state at low temperatures and a peculiar linear negative temperature coefficient of resistance in the high-temperature regime with a predominant electron-phonon scattering mechanism. Analysis of room temperature powder X-ray diffraction data reveals a highly ordered L21 structure and reduction of antisite disorder, upon Mn substitution. The temperature-dependent magnetization measurements reveal distinct features indicative of weak anisotropy in the system. Isothermal magnetization measured as a function of the applied field helps identify the unique magnetic ground state of the half-metallic Fe2Ti1-xMnxSn composition as a ferromagnet with a wandering axis, that distinctively orients in the direction of applied field. Our findings thus provide a new perspective for studying the mechanism of half-metallicity and associated magnetic order in Heuslers.
@@ I Introduction
Since its discovery in 1903[1], Heusler alloys with general formula X2YZ (where X and Y are transition metals and Z is the main group element), have constantly drawn attention of the researchers due to its remarkable properties, such as magnetic shape memory effect[2], near - room temperature magnetocaloric effect[3], half-metallic ferromagnetism[4, 5], spin-gapless semi-conductivity, unconventional superconductivity[6] and anomalous Hall effect, and so on. These compositions crystallize in the L21 structure (space group: Fm-3m) consisting of 4 interpenetrating FCC sublattices, where X atom occupies the (0.25,0.25,0.25) and (0.75,0.75,0.75) sublattice sites, whereas the Y and Z atoms occupy the (0.5,0.5,0.5) and (0,0,0) sublattice sites, respectively. Theoretical studies on Heusler alloys reveal an interesting relationship between the total spin moment per unit cell (MT) and the number of valence electrons count (VEC) per formula unit (Z). Also known as Slater-Pauling rule[7], this relationship states quantitative value of MT as equal to (Z - 24), and seems to provide a handle to modulate the physical properties of Heuslers through suitable chemical substitutions.
@@
Ideally, for compositions where VEC equals 24, the Slater-Pauling rule suggests a non-magnetic ground state. Examples include the Fe2VAl, Fe2VGa[8, 9] compositions. However, experimental studies of the magnetic properties show signatures of weak magnetism[10, 11], the reason being presence of anti-site and quenched disorders[12]. Heusler alloys with VEC 24 are also considered for applications such as magnetic information storage where they could play the role of a thin non-magnetic buffer layer sandwiched between two ferromagnetic layers of other Heusler compositions, acting as a spin polarizer in a recording head while providing the very compatible inter-layer coupling. Furthermore, a significant decrease in the resistivity, magnetization, and magnetoresistance values has been reported When other transition metal atom are substituted at X or Y sites, that can lead to a state of high spin-polarization[13] at Fermi level. A narrow gap/ pseudogap forms in the vicinity of the Fermi level with transport measurements reflecting a semiconducting nature and heavy effective mass of charge carriers. These properties make them suitable for thermoelectric applications.
@@
Another notable example is the Fe2TiSn composition, which conforms to the VEC = 24 prototype. The electronic structure calculation indicates Fe2TiSn to be a non-magnetic semimetal with a pseudogap in the density of state at Fermi level. Based on specific heat measurements, it has been speculated to be a heavy fermion metal with quasi-particle mass 40 me[14]. The substitution of Ni at Fe sites forms a system with isolated spins leading to a Kondo-like interaction between localized spins and conduction electrons [15]. On the contrary, Co substitution leads to a typical spin fluctuation scenario. Optical conductivity and infrared studies provide evidence of pseudogap in the density of states and, the mass enhancement at low temperature to be Schottky anomaly from magnetic clusters resulting from antisite disorder between Fe and Ti sites. In contrast, Sayan et al. have shown that the substitution of a small percentage of Ti at the Fe site reduces antisite disorder and resistivity and specific heat measurements suggest the presence of spin fluctuation, which decreases with decreasing disorder. It is to be noted here that in contrast to Fe and Co, bulk Ni has a sharp peak at Ef in one of the spin sub-bands[16]. This sudden increase of DOS at EF may lead to enhanced electron-electron interaction (EEI) leading to Kondo-like spin scattering at low temperature. Furthermore, the non-magnetic semi-metallic band structure of Fe2TiSn demonstrates a low density of states at the Fermi level, rendering the electronic structure and magnetic properties highly sensitive to doping or elemental substitution and with proper substitution at Ti and Sn sites, the system can be driven to half-metallicity[17, 18].
@@
In this work, we report the detailed investigation of the structural, magnetic, and transport properties of Mn-substituted Fe2TiSn compositions. X-ray diffraction (XRD) results show the formation of an ordered L21 structure with highly reduced antisite disorder. Temperature and field-dependent magnetization measurements indicate the presence of anisotropy in the system, due to the strong local magnetic moment of Mn atoms. The low-temperature upturn in the electrical resistivity is due to the combined effect of electron-electron correlation and strong Kondo-like spin scattering. The (T) above 100 K can be best explained by electron-phonon and electron-magnon scattering, and the contribution of electron-magnon scattering is comparatively less, which indicates the presence of spin polarization of density of state in the prepared composition. We have also performed the electronic structure calculations to better understand the ground state properties.
@@ II Experimental and Computational details
Polycrystalline Fe2Ti1-xMnxSn compositions were prepared by melting stoichiometric amounts of high purity (99.99%) elements such as, Fe, Mn, Ti, and Sn using an electric arc furnace by placing these elements on a water-cooled Cu hearth and under flowing Ar atmosphere. The homogeneous ingots thus formed were vacuum sealed in the quartz tubes and annealed for 72 hours at 800∘C followed by quenching in ice water. The weight loss during the whole process was 1. The energy dispersive X-ray spectroscopy (EDS), equipped with a field emission scanning electron microscope was performed using Model FE-SEM Supra 55 (Carl Zeiss, Germany), to ensure the target composition after melting was close to the starting stoichiometry. The atomic percentages of the elements are presented in Table1 and the elemental mapping for = 0.25 is shown in Fig.1. The crystallographic phase purity of the samples checked using a laboratory X-ray source, on Bruker D8 Advance diffractometer having Cu target ( = 1.5406 ), confirms the formation of the Heusler phase. Magnetization as a function of temperature and magnetic field is measured using a Quantum Design SQUID magnetometer. A standard four-probe method with a homemade resistivity setup equipped with a superconducting magnet system was used for the electrical resistivity and magnetoresistance measurements.
Density-functional theory (DFT) calculations are performed within the Quantum Espresso package [19, 20] where ultrasoft pseudopotentials [21] and a plane-wave basis set are employed. Generalized-gradient approximation (GGA) with Perdew–Burke–Ernzerhof (PBE) functional form is used for the calculation of exchange-correlation energy [22]. For Sn, Ti, Mn and Fe, 455, 3343, 3434 and 3433 are treated as valence states respectively. The kinetic energy cutoff for plane-wave functions of 95 Ry and the charge density of 1140 Ry are considered in our calculations. We start with the simple cubic crystal structure (space group: , no: 225) of Fe2TiSn with lattice constant = 6.059 Å and fully optimize it after replacing one of the Ti atoms with Mn (25 % doping). Energy convergence criterion of 5 10-7 Ry and 666 uniform -mesh and Marzari-Vanderbilt smearing are considered in our calculations. The optimization leads to small changes ( 1.0 %) in the lattice constants. These spin-polarized calculations are performed starting with aligning all eight Fe atoms ferromagnetically with large moments while the one Mn atom is aligned antiferromagnetically to the Fe atoms in the total 16–atom unit cell. However, after the convergence, the negligibly small Fe and large Mn magnetic moments are aligned ferromagnetically. We discuss these results in connection with the magnetic properties of the compositions, at a later stage.
@@ III Results and Discussion
The density of states (DOS) scenario of the Fe2Ti1-xMnxSn system is obtained using a DFT-only approach. The atom-projected spin-polarized DOS of Fe, Mn, Ti - and Sn states for this case are shown in Fig. 2. Many insightful observations can be made from this plot. First, the system shows the half-metallic character of the DOS with mainly Fe and Mn states in the spin-up channel, while there is a band gap in the spin-down channel at the Fermi level. Though Ti states also contribute below the Fermi level, the dominant contributors remain to be the states of Fe and Mn atoms, and the magnetic properties arise from the interactions among all three transition metal atoms. Second, this interaction can safely be assumed to be itinerant as the system is metallic. Third, the magnetic moment on the Mn site is large, while Fe and Ti atoms carry negligible moment while both, Mn & Fe have finite contribution to the DOS at the Fermi level. Recall that we started with a scenario of antiferromagnetic coupling between single Mn and eight (ferromagnetically aligned) Fe atoms. Such a system converges to a ferromagnetic coupling among all the magnetic atoms, with negligibly small magnetic moments on Fe ( 0.016 ) and Ti ( 0.052 ) atoms, and a relatively high moment on Mn ( 3.161 ) atoms. It appears as if Mn atoms are the ones with localized moments that are interacting through the cloud of conduction electrons formed by Fe and Ti atoms.
@@
With the intent of finding a half-metallic Heusler, we begin the experimental investigation into the Mn–substituted Fe2TiSn compounds. The X-ray diffraction (XRD) patterns of Fe2Ti1-xMnxSn ( = 0.05, 0.1, 0.25, and 0.3) are shown in Fig.3(a). The presence of strong (111), (200), and (220) peaks confirm the highly ordered cubic L21 crystal structure in all the prepared compositions. However, the solubility limit of Mn in Fe2TiSn is limited to 30 as additional peaks related to the secondary phase begin to appear for compositions with 0.4. Contrary to the general expectation of finding an increase in the antisite disorder upon carrying out substitutions in an intermetallic system, the present compositions show a relative increase in the intensity of the (111) and (200) super-lattice peaks, indicating a reduction of the overall disorder in comparison to pristine Fe2TiSn. Rietveld refinement of the XRD profiles, implemented through the FullProf suite[23], yield values for lattice parameters that follow Vegard’s law[24] indicating the desired substitutions at the targeted lattice sites have indeed been achieved.
With the prediction of half-metallic DOS at the Fermi level and a highly ordered crystal structure, it becomes pertinent to study the electrical resistivity of the Mn–substituted Fe2TiSn compounds. Fig.4 presents the (T) measured for all the compositions in the temperature range of 5–-320 K and with a constant magnetic field of 0 and 5 T. For low Mn–concentration of = 0.05 and 0.1, (T) decreases with decreasing temperature, displaying the expected metallic character. However, for the half-metallic composition of = 0.25 (and also for = 0.3), (T) changes dramatically with a negative temperature coefficient of resistance (TCR) in the given temperature range. The overall magnitude of resistivity also increases as changes from 0.05 to 0.3, while in principle, overall valence electron count increases with the systematic replacement of Ti with Mn takes place. A careful analysis of (T) in the different temperature segments hence becomes necessary to bring to the fore the underlying charge scattering mechanisms.
Firstly, the presence of a narrow gap in one of the spin sub-bands that has its origin in a half-metallic band structure can give rise to an increased magnitude of resistivity due to the absence of the spin-flip level in the electron-magnon scattering process[25, 26, 27]. A negative TCR in = 0.25 (and 0.3) is also not an unusual trend in Heuslers and many quaternary and half-Heusler compounds are known to show such dependence[28, 29, 30]. However, the interesting aspect here is the near linear (T) between 100 to 300 K, rather than a thermally activated power law or an exponentially decaying function. A good fit to the data in this temperature region is obtained through a combination of the highly dominant electron-phonon scattering process (T linear term) along with a small contribution of electron-magnon scattering (T2 term), as described by the equation, and displayed in Fig.5. Since the scattering is a spin-flip process, the whole spin wave needs to consist of both the spin-up and spin-down sub-bands. The significantly low value of as compared to the denotes the absence of one of the spin density of state through opening of the gap at the Fermi level, leading to the formation of the spin-polarized band structure, and agreeing with our electronic structure calculations.
10-4 | 10-6 | 10-5 | 10-14 | 10-7 | 10-11 | |||||
@@
0.25 | 9.905 | 9.171 | 8.083 | 10.747 | 1.786 | 1.388 | 6.554 | 19.91 | 9.95 | 5.42 |
0.30 | 10.6 | 9.965 | 4.029 | 8.318 | 1.372 | 1.346 | 84.906 | 24.386 | 1.179 | 25.33 |
In the same temperature range, the compositions with low Mn concentration of = 0.05, 0.1 display the expected metallic behavior, with its characteristic broad hump–feature. On the high-temperature side of the hump–feature lies the semi-metallic nature of Fe2TiSn, while the low-temperature side is the typical metallic conduction[31]. The reason for the broad hump–feature itself, is the short-range magnetic clusters that are known to form even in pristine Fe2TiSn due to the sizable antisite disorder between Fe & Ti atoms, as already shown in literature[17]. In the compositions containing low Mn percentages, the peak of the hump–feature is shifted to a slightly lower temperature. The sizeable effect of antisite disorder is also seen at the low temperature, wherein (T) experiences an upturn for = 0.05 and 0.1. Recent studies have shown that the application of a constant external magnetic field overcomes this upturn and signifies the presence of weak localization in Fe2TiSn[17] and the same scenario applies to these low Mn-substituted compositions.
On the other hand for compositions with = 0.25, 0.3, the negative TCR in continues until a cusp-like feature occurs at temperatures below 20 K. This feature gets suppressed upon application of a constant external magnetic field, without any effect on the high-temperature linear resistivity. The data in the low-temperature region does not fit the simple description of weak localization, as was the case for = 0.05, 0.1 compositions. Rather, a good fit to the data below 20 K and for H = 0 & 5 T is obtained using the description,
, which encompasses the electron-electron () interaction that follows a dependence, the Kondo-like spin scattering with the T term, and the electron-phonon () scattering that goes as . The fit to the data is presented in Fig.5, and the fitting parameters are given in Table.2. The magnitude of is one order higher than , indicating the dominance of Kondo-like scattering in these high Mn-content compositions. Our electronic structure calculations already point towards a scenario where localized moments reside at the Mn-sites that interact through the cloud of conduction electrons formed by Fe and Ti atoms.
@@
Alternatively, the magnetic field dependence of electrical resistance at fixed temperatures also presents the same scenario as that of (T). Fig.6 shows the isothermal magnetoresistance (MR) curves measured at different temperatures ranging from 5 KT300 K and for all four compositions. Using the standard definition of MR as, 100, we find that all the Mn–substituted compositions show the isotropic negative MR in the entire temperature range. Due to the dominance of electron-phonon scattering in the high-temperature regime, the MR curves exhibit linearity with changing magnetic fields. However, at low temperatures, a slight deviation from linearity is observed. Besides, in = 0.25, 0.3, a sharp increase in the magnitude of the MR is observed at temperatures below 20 K, where the cusp-like feature was observed in (T). A power law fitting to the low -temperature MR data provides an exponent value of 0.471, which is as per the Altshuler-Aronov model that predicts an electron-electron correlation dominated MR to have a H1/2 dependence[32].
@@
Having understood the electronic DOS in the Fe2Ti1-xMnxSn compositions and half-metallic ground state in 0.25, we now discuss the experimental study of its magnetic properties. The temperature-dependent magnetization, M(T), plotted in Fig.7. According to the previous studies, an inherent Fe/Ti antisite disorder in the otherwise non-magnetic Fe2TiSn induces the formation of weak magnetic clusters[33]. These interactions are identified by a step-like feature in the high-temperature region, denoted here as TS, in the M(T) curve. This step-like feature shifts towards the high-temperature side with rising Mn substitution. Additionally, at low–temperatures an abrupt rise in magnetization is observed marking the transition to a magnetically ordered state. This temperature is identified as TC and its value increases with rising Mn content. The value of TC and TS are listed in Table.1.

### v2

# Half-metallicity and wandering axis ferromagnetism in Fe2Ti1-xMnxSn (0 0.3) Heusler Alloys

@@ Abstract
We investigate the effect of Mn substitution in Fe2Ti1-xMnxSn on electronic structure and magnetic and electrical transport properties. The spin-polarized density of states calculations using density-functional theory (DFT) yields a half-metallic ground state in Mn-rich compositions. Localized magnetic moments at Mn sites interacting through the cloud of conduction electrons formed by Fe and Ti atoms are also predicted. Electrical resistivity and magneto-transport measurements reveal a Kondo-like ground state at low temperatures and a peculiar linear negative temperature coefficient of resistance in the high-temperature regime with a predominant electron-phonon scattering mechanism. Analysis of room temperature powder X-ray diffraction data reveals a highly ordered L21 structure and reduction of antisite disorder upon Mn substitution. The temperature-dependent magnetization measurements reveal distinct features indicative of weak anisotropy in the system. Isothermal magnetization measured as a function of the applied field helps identify the unique magnetic ground state of the half-metallic Fe2Ti1-xMnxSn composition as a ferromagnet with a wandering axis that distinctively orients in the direction of the applied magnetic field. The measurement of X-ray absorption fine structure (XAFS) reveals that the random anisotropy arises due to the local lattice distortion around Mn atoms in the prepared compositions. Our findings thus provide a new perspective for studying the mechanism of half-metallicity and associated magnetic order in Heuslers.
@@ I Introduction
Since its discovery in 1903[1], Heusler alloys with general formula X2YZ (where X and Y are transition metals and Z is the main group element) have constantly drawn the attention of the researchers due to their remarkable properties, such as magnetic shape memory effect[2], near - room temperature magnetocaloric effect[3], half-metallic ferromagnetism[4, 5], spin-gapless semi-conductivity, unconventional superconductivity[6] and anomalous Hall effect, and so on. These compositions crystallize in the L21 structure (space group: Fm-3m) consisting of 4 interpenetrating FCC sublattices, where X atom occupies the (0.25,0.25,0.25) and (0.75,0.75,0.75) sublattice sites, whereas the Y and Z atoms occupy the (0.5,0.5,0.5) and (0,0,0) sublattice sites, respectively. Theoretical studies on Heusler alloys reveal an interesting relationship between the total spin moment per unit cell (MT) and the number of valence electrons count (VEC) per formula unit (Z). Also known as Slater-Pauling rule[7], this relationship states quantitative value of MT as equal to (Z - 24), and seems to provide a handle to modulate the physical properties of Heuslers through suitable chemical substitutions.
@@
Ideally, for compositions where VEC equals 24, the Slater-Pauling rule suggests a non-magnetic ground state. Examples include the Fe2VAl, Fe2VGa[8, 9] compositions. However, measurement of the magnetic properties of these compositions show signatures of weak magnetism[10, 11], the reason being the presence of anti-site and quenched disorders[12]. Heusler alloys with VEC 24 are also considered for applications such as magnetic information storage, where they could play the role of a thin non-magnetic buffer layer sandwiched between two ferromagnetic layers of other Heusler compositions, acting as a spin polarizer in a recording head while providing the very compatible inter-layer coupling. Furthermore, a significant decrease in the resistivity, magnetization, and magnetoresistance values has been reported when other transition metal atoms are substituted at X or Y sites, which can lead to a state of high spin-polarization[13] at Fermi level. A narrow gap/pseudogap forms near the Fermi level, with transport measurements reflecting a semiconducting nature and heavy effective mass of the charge carriers, making them suitable for thermoelectric applications.
@@
Another notable example is the Fe2TiSn composition, which conforms to the VEC = 24 prototype. The electronic structure calculation indicates Fe2TiSn to be a non-magnetic semimetal with a pseudogap in the density of state at the Fermi level. Based on specific heat measurements, it has been speculated to be a heavy fermion metal with quasi-particle mass 40 me[14]. The substitution of Ni at the Fe site forms a system with isolated spins, leading to a Kondo-like interaction between localized spins and conduction electrons [15]. On the contrary, Co substitution leads to a typical spin fluctuation scenario. Optical conductivity and infrared studies of Fe2TiSn provide evidence of pseudogap in the density of states and the mass enhancement at low temperatures to be Schottky anomaly from magnetic clusters resulting from antisite disorder between Fe and Ti sites. It is to be noted here that in contrast to Fe and Co, bulk Ni has a sharp peak at Ef in one of the spin sub-bands[16]. This sudden increase of DOS at EF may lead to enhanced electron-electron interaction (EEI), leading to Kondo-like spin scattering at low temperatures. Furthermore, the non-magnetic semi-metallic band structure of Fe2TiSn demonstrates a low density of states at the Fermi level, rendering the electronic structure and magnetic properties highly sensitive to doping or elemental substitution and with proper substitution at Ti and Sn sites, the system can be driven to half-metallicity[17, 18].
@@
In this work, we report the detailed investigation of the structural, magnetic, and transport properties of Mn-substituted Fe2TiSn compositions. X-ray diffraction (XRD) results show the formation of an ordered L21 structure with highly reduced antisite disorder. Temperature and field-dependent magnetization measurements indicate the presence of anisotropy in the system due to the strong local magnetic moment of Mn atoms. The low-temperature upturn in the electrical resistivity is due to the combined effect of electron-electron correlation and strong Kondo-like spin scattering. The (T) above 100 K can be best explained by electron-phonon and electron-magnon scattering, and the contribution of electron-magnon scattering is comparatively less, which indicates the presence of spin polarization of density of state in the prepared composition. We have also performed the electronic structure calculations to better understand the ground state properties.
@@ II Experimental and Computational details
Polycrystalline Fe2Ti1-xMnxSn compositions were prepared by melting stoichiometric amounts of high purity (99.99%) elements such as Fe, Mn, Ti, and Sn using an electric arc furnace by placing these elements on a water-cooled Cu hearth and under flowing Ar atmosphere. The homogeneous ingots thus formed were vacuum sealed in the quartz tubes and annealed for 72 hours at 800∘C followed by quenching in ice water. The weight loss during the whole process was 1. The energy dispersive X-ray spectroscopy (EDS), equipped with a field emission scanning electron microscope, was performed using Model FE-SEM Supra 55 (Carl Zeiss, Germany) to ensure the target composition after melting was close to the starting stoichiometry. The atomic percentages of the elements are presented in Table-1, and the elemental mapping for = 0.25 is shown in Fig.1. The crystallographic phase purity of the samples was checked using a laboratory X-ray source on Bruker D8 Advance diffractometer having Cu target ( = 1.5406 Å), confirming the formation of the Heusler phase. Magnetization as a function of temperature and magnetic field was measured using a Quantum Design SQUID magnetometer. A standard four-probe method with a homemade resistivity setup equipped with a superconducting magnet system was used for the electrical resistivity and magnetoresistance measurements. The Extended X-ray absorption fine structure (EXAFS) measurement was carried out at P65 beamline, PETRA-III DESY, Hamburg, Germany at Fe (7112 eV) and Ti (4966 eV) K edges at 6 K. Gas-filled ionization chambers were used to simultaneously record the incident () and transmitted () photon energies. The data was collected over various scans to reduce the statistical noise. Subsequently, the EXAFS signal was extracted and analyzed in the 3 to 14 Å-1 range using Demeter suite[19].
Density-functional theory (DFT) calculations are performed within the Quantum Espresso package [20, 21] where ultrasoft pseudopotentials [22] and a plane-wave basis set are employed. Generalized-gradient approximation (GGA) with Perdew–Burke–Ernzerhof (PBE) functional form is used for the calculation of exchange-correlation energy [23]. For Sn, Ti, Mn and Fe, 455, 3343, 3434 and 3433 are treated as valence states respectively. The kinetic energy cutoff for plane-wave functions of 95 Ry and the charge density of 1140 Ry are considered in our calculations. We start with the simple cubic crystal structure (space group: , no: 225) of Fe2TiSn with lattice constant = 6.059 Å and fully optimize it after replacing one of the Ti atoms with Mn (25 % doping). Energy convergence criterion of 5 10-7 Ry and 666 uniform -mesh and Marzari-Vanderbilt smearing are considered in our calculations. The optimization leads to small changes ( 1.0 %) in the lattice constants. These spin-polarized calculations start with aligning all eight Fe atoms ferromagnetically with large moments while the one Mn atom is aligned antiferromagnetically to the Fe atoms in the total 16–atom unit cell. However, after the convergence, the negligibly small Fe and large Mn magnetic moments are aligned ferromagnetically. We discuss these results in connection with the magnetic properties of the compositions at a later stage.
@@ III Results and Discussion
The density of states (DOS) scenario of the Fe2Ti1-xMnxSn system is obtained using a DFT-only approach. For this case, the atom-projected spin-polarized DOS of Fe, Mn, Ti - and Sn states are shown in Fig.2. Many insightful observations can be made from this plot. First, the system shows the half-metallic character of the DOS with mainly Fe and Mn states in the spin-up channel, while there is a band gap in the spin-down channel at the Fermi level. Though Ti states also contribute below the Fermi level, the dominant contributors remain the states of Fe and Mn atoms, and the magnetic properties arise from the interactions among all three transition metal atoms. Second, this interaction can be considered itinerant as the system is metallic. Third, the magnetic moment on the Mn site is large, while Fe and Ti atoms carry negligible moments, and both Mn & Fe have finite contributions to the DOS at the Fermi level. Recall that we started with a scenario of antiferromagnetic coupling between single Mn and eight (ferromagnetically aligned) Fe atoms. Such a system converges to a ferromagnetic coupling among all the magnetic atoms, with negligibly small magnetic moments on Fe ( 0.016 ) and Ti ( 0.052 ) atoms and a relatively high moment on Mn ( 3.161 ) atoms. It appears as if Mn atoms are the ones with localized moments that are interacting through the cloud of conduction electrons formed by Fe and Ti atoms.
@@
With the intent of finding a half-metallic Heusler, we begin the experimental investigation into the Mn–substituted Fe2TiSn compounds. The X-ray diffraction (XRD) patterns of Fe2Ti1-xMnxSn ( = 0.05, 0.1, 0.25, and 0.3) are shown in Fig.3(a). The presence of strong (111), (200), and (220) peaks confirm the highly ordered cubic L21 crystal structure in all the prepared compositions. However, the solubility limit of Mn in Fe2TiSn is limited to 30 as additional peaks related to the secondary phase begin to appear for compositions with 0.4. Contrary to the general expectation of finding an increase in the antisite disorder upon carrying out substitutions in an intermetallic system, the present compositions show a relative increase in the intensity of the (111) and (200) super-lattice peaks, indicating a reduction of the overall disorder in comparison to pristine Fe2TiSn. Rietveld refinement of the XRD profiles, implemented through the FullProf suite[24], yield values for lattice parameters that follow Vegard’s law[25], indicating the desired substitutions at the targeted lattice sites have indeed been achieved.
With the prediction of half-metallic DOS at the Fermi level and a highly ordered crystal structure, studying the electrical resistivity of the Mn–substituted Fe2TiSn compounds becomes pertinent. It has been previously observed that the electrical resistivity, , of Fe2TiSn, shows a bad metallic behavior with a low-temperature upturn and minimum at 50 K. This low-temperature upturn in is due to weak localization caused by anti-site disorder between the Fe and Ti sites. Fig.4 shows the data for Fe2Ti1-xMnxSn ( = 0.05, 0.10, 0.25, 0.30) for a temperature range of 5–300 K and at applied magnetic fields of 0 T and 5 T. For brevity, each composition’s is normalized to value. The general features of curve of = 0.05 sample match that of Fe2TiSn[26]. However, when viewed from the high-temperature side of the curve for = 0.10 sample shows a negative temperature coefficient of resistance (TCR) down to 200 K, followed by positive TCR at low temperatures. On the other hand, = 0.25 and 0.30 compositions exhibit complete negative TCR throughout the measurement range. A closer inspection reveals that the of = 0.25 and 0.30 shows a weak slope change at low temperature (indicated by arrow), below which increases at a faster rate with decreasing temperature.
10-2 | 10-2 | 10-9 | 10-3 | 10-7 | ||||||
@@
0.25 | 1.389 | 1.292 | 0.67 | 1.20 | 1.89 | 1.7 | 1.239 | 0.73 | -1.18 | -5.909 |
0.30 | 1.376 | 1.297 | 1.01 | 1.24 | 2.5 | 1.9 | 0.397 | 0.135 | -1.29 | -2.42 |
The low-temperature data is analyzed by considering the different scattering mechanisms that may contribute to the observed behavior. A low- upturn in resistivity can result from various mechanisms, including electron-electron interaction (EEI), weak localization (WL), and Kondo-like transport. EEI arises from electron-electron elastic scattering due to the Coulomb interaction between conduction electrons, WL arises from disorder-mediated coherent backscattering of charge carriers, and Kondo-like transport arises from the interaction between localized magnetic moment associated with the magnetic impurities and mobile conduction electrons[27, 28]. In Mn substituted Fe2TiSn, a reduction in atomic disorder and/or magnetic order destroys the WL, suggesting that the low-T upturn of in compositions with higher Mn concentration is not caused by WL[26, 29]. Furthermore, we fit the data separately with EEI and Kondo effect. As the result indicates, neither model fits the experimental data satisfactorily. Thus, combining both mechanisms likely contributes to the observed low-T upturn in resistivity. Besides, the contribution from electron-phonon scattering should also be added to the form to represent the resistivity[30]. Thus, the low-T resistivity data at H = 0 T and 5 T can be described by the expression
, where, is residual resistivity, and, , and represent coefficients for EEI, Kondo-like transport and electron–phonon scattering respectively. The fitting result is shown in Fig.5, and the fitting parameters for different Mn substitution and applied magnetic field are depicted in Table-2. Both and increase with increasing . The value of for a given composition is higher at 0 T compared to 5 T, highlighting the significant role of magnetic scattering of charge carriers. Whereas not only the value of is higher at 5 T as compared to that of 0 T, but it also has a positive value. However, if the effect of both Kondo and EEI are simultaneously taken into account, may end up being positive[31]. Generally, when EEI alone contributes to the , the value extracted from fitting the data yields a negative value.
Next, we analyze the behavior of of = 0.25 ( and 0.30) composition in the high-T region. While the negative TCR has been observed in Heusler compounds[32, 33, 34, 35, 36], the notable feature here is the linear between 100 K and 300 K, which is unusual compared to the common semiconductor and disordered metallic systems. In semiconductors, electron transport exhibits a thermally activated behavior of data where vs T plot is upward convex. On the contrary, disordered metallic systems show a convex downward behavior of data. Further, reports on the high-T resistivity data of Fe2TiSn show a negative TCR from 300 K to 800 K, which is attributed to the interband transition of conduction electrons through a gap at Fermi level[37]. With electron doping in pristine Fe2TiSn, this negative TCR persists below 300 K[15] and may also result in half-metallicity[17, 18]. A similar signature is observed in the present Fe2Ti1-xMnxSn compositions. In the case of half-metals, the high-temperature electrical resistivity can be expressed by the combined contribution from electron-magnon and electron-phonon scattering. The contribution of electron-magnon scattering towards resistivity follows T2 dependence and the contribution of the scattering of conduction electron by phonons is linear-T at high temperatures. Thus, we have fitted the resistivity data in the high-T region using the equation:
, where and represent the coefficients of electron-phonon and electron-magnon scattering, respectively. The fitting results are shown in Fig.5. Two observations can be made from the values of the and (given in Table-2), both the coefficients are negative and the magnitude of is very small compared to the . The negative value of the coefficients is expected as increases with a decrease in temperature[32]. Additionally, since electron-magnon scattering is a spin-flip process, the lower value of as compared to the denotes the presence of spin-polarised DOS at Fermi level in = 0.25 (and 0.30). Our theoretical calculations for = 0.25 also support this observation. Thus, the combined results of theoretical calculations and electrical resistivity indicate the presence of half metallicity in Mn-substituted Fe2TiSn.
@@
Alternatively, the magnetic field dependence of electrical resistance at fixed temperatures also presents the same scenario as that of (T). Fig.6 shows the isothermal magnetoresistance (MR) curves measured at different temperatures ranging from 5 KT300 K and for all four compositions. Using the standard definition of MR as, 100, we find that all the Mn–substituted compositions show the isotropic negative MR in the entire temperature range. Generally, the field-dependent resistivity, , depends on the combined sum of Lorentz force and contribution from different scattering mechanisms. The Lorentz force constitutes the classical MR, which is inherently positive. Meanwhile, in the studied compositions, the high-T electrical resistivity is mainly governed by electron-phonon and electron-magnon scattering, where the former is nearly independent of the applied magnetic field. The combined effect of these two processes likely leads to the observed small yet negative MR at high temperatures[38].
Moreover, the increase in the magnitude of the low-T MR with increasing Mn content suggests the presence of magnetic ordering and/or spin-dependent scattering. Additionally, compositions showing signatures of half-metallicity, i.e. = 0.25 (and 0.30), exhibit the dominance of EEI and the Kondo effect in low-T resistivity. Notably, the EEI effect leads to a positive MR due to Zeeman splitting and orbital effects, with its magnitude varying as the square root of the applied magnetic field[27, 39, 40, 38, 41]. This has been reflected as the increment of value with an applied magnetic field of 5 T for a given value of (see Table-2). Conversely, applying the magnetic field suppresses the fluctuation in the localized magnetic moment and spin-dependent scattering, leading to a negative MR in a Kondo system[42]. Thus, for = 0.25 (and 0.30), the negative MR supports that the Kondo effect is the dominant mechanism governing the low-T electrical resistivity. This also explains the higher value of as compared to the obtained from the fitting of the low-T resistivity data.
@@
Having understood the electronic DOS in the Fe2Ti1-xMnxSn compositions and half-metallic ground state in 0.25, we now discuss the experimental study of its magnetic properties. The temperature-dependent magnetization, , measured at 0.01 T, is plotted in Fig.7. According to the previous studies, an inherent Fe/Ti antisite disorder in the otherwise non-magnetic Fe2TiSn induces the formation of weak magnetic clusters[26]. These interactions are identified by a step-like feature in the high-temperature region, denoted here as TS, in the M(T) curve. This step-like feature shifts towards the high-temperature side with rising Mn substitution. Additionally, an abrupt rise in magnetization is observed at low– temperatures, marking the transition to a magnetically ordered state. This temperature is identified as TC, and its value increases with rising Mn content. The values of TC and TS are listed in the Table-1.